# Classically-embedded split Cayley hexagons rule three-qubit contextuality with three-element contexts

### v1

@@ Abstract
As it is well known, split Cayley hexagons of order two live in the three-qubit symplectic polar space in two non-isomorphic embeddings, called classical and skew. Although neither of the two embeddings yields observable-based contextual configurations of their own, classically-embedded copies are found to fully rule contextuality properties of the most prominent three-qubit contextual configurations in the following sense: each set of unsatisfiable contexts of such a contextual configuration is isomorphic to the set of lines that certain classically-embedded hexagon shares with this particular configuration. In particular, for a doily this shared set comprises three pairwise disjoint lines belonging to a grid of the doily, for an elliptic quadric the corresponding set features nine mutually disjoint lines forming a (Desarguesian) spread on the quadric, for a hyperbolic quadric the set entails 21 lines that are in bijection with the edges of the Heawood graph and, finally, for the configuration that consists of all the 315 contexts of the space its 63 unsatisfiable ones cover an entire hexagon. A particular illustration of this encoding is provided by the line-complement of a skew-embedded hexagon; its 24 unsatisfiable contexts correspond exactly to those 24 lines in which a particular classical copy of the hexagon differs from the considered skew-embedded one. In connection with the last-mentioned case we also conducted some experimental tests on a Noisy Intermediate Scale Quantum (NISQ) computer to validate our theoretical findings.
@@ 1 Introduction
Finite geometry is a branch of mathematics that deals with geometries made of a finite number of points, lines and/or linear spaces of higher dimensions. This perspective of working with spaces that contain only a finite number of geometric elements is rather counter-intuitive and far from the intuition provided, for example, by Euclidean geometry as it lacks concepts like smoothness, differentiability, distance etc. Gino Fano [fano] was one of the first geometers to formalize this idea and his name is now associated with the smallest finite projective plane, i. e. the Fano plane (Figure 1).
Figure 1: The Fano plane depicted in its standard rendering. This self-dual geometry comprises seven points (bullets) and seven lines (six straight segments and a circle), with three points per line and, dually, three lines through a point, being equivalent to the projective plane over the two-elements field $\mathbb{F}_{2}=\{0,1\}$.
Over the past 20 years (see, for example, [sp06], [hs08], [slp12], [lps13] and [ls17]), finite geometry has been introduced into the field of quantum information and mathematical physics to mainly model and analyse the commutation relations within the $n$-qubit Pauli group, $\mathcal{P}_{n}$, defined by
@@
where $A_{1}A_{2}\dots A_{n}\equiv A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}$. For example, for $n=3$, if one ignores the global phase of each operator, a maximum set of mutually commuting three-qubit Pauli operators (disregarding the identity) forms a Fano plane, as shown with the example provided by Figure 2. Two points are colinear if and only if the matrix product of the observables labelling them commute.
@@
In what follows, a (quantum) context will be a set of mutually commuting observables such that their product is ±$\pm I^{\otimes n}$. For instance, the seven observables of Figure 2 form a (negative) three-qubit context. A contextual configuration will be an arrangement of observables made of contexts such that there is no Non-Contextual Hidden Variable (NCHV) model that can reproduce the outcomes predicted by the rules of Quantum Mechanics (QM).
@@
Let us consider, for example, the configuration portrayed in Figure 3, known as the Mermin pentagram. Each node of the configuration is a three-qubit Pauli operator whose eigenvalues are $+1$ and $-1$. Each line of the configuration represents a context. The doubled line is the unique context with a negative sign, i. e. the product of the observables on this context is $-I^{\otimes 3}$. David Mermin introduced this configuration in [mermin] as an alternative proof of the famous Kochen-Specker (KS) Theorem [kochen1990problem]. Recall that the Kochen-Specker Theorem is a no-go result that proves the non-existence of an NCHV model, i. e. it proves that a Hidden Variable (HV) model that would reproduce the outcomes predicted by QM has to be context-dependent. Consider our Mermin pentagram. An HV model that reproduces predictions of QM for this set of operators should satisfy the following constraints:
@@
Observable-based proof of the KS Theorem can already be built from two-qubit Pauli observables, like in the Mermin-Peres magic square, see Section 2 and  [peres1991two, mermin]. For more information on the two-qubit case, the reader is referred to consult [DHGMS22, HS17], as well as [budroni2022kochen] for a recent survey on contextuality.
@@
In this paper we will focus solely on the three-qubit case ($n=3$) to demonstrate in detail how a specific arrangement of three-qubit contexts isomorphic to the smallest split Cayley hexagon underpins contextuality properties of a whole class of distinguished aggregates of three-qubit observables. This hexagon made its debut in physics as early as 2008 in [lsv], where the authors employed its particular three-qubit subgeometry to model the $E_{7}$-symmetric black-hole entropy formula in string theory. Later [psh], it was pointed out that the order of the automorphism group of the hexagon, 12 096, coincides with the total number of three-qubit Mermin pentagrams. At about the same time [sppl], it was already noticed that some contextual three-qubit configurations can be uniquely extended to geometric hyperplanes of the hexagon.
@@
It is however only very recently [holweck2022three] that the fact that the hexagon embeds in the three-qubit space in two different ways has been properly taken into account and recognized as having deep physical meaning, leading to our present study.
@@
The paper is organized as follow. In Section 2 we introduce the notion of the symplectic polar space of rank $n$ and order two, $\mathcal{W}(2n-1,2)$, which is the geometric framework for the commutation relations within the $n$-qubit generalized Pauli group. Next, we define the most prominent subgeometries of this space, quadrics, and show that the particular space $\mathcal{W}(5,2)$ for three-qubits is endowed with another remarkable kind of subgeometry, namely the split Cayley hexagon of order two, which occurs in this space in two non-isomorphic embeddings. Then we introduce the notion of the degree of contextuality of a quantum configuration, that measures how far a contextual configuration of observables is from supporting an NCHV model, and explicitly illustrate this notion on the example of the smallest symplectic space $\mathcal{W}(3,2)$, the doily, and one of its hyperbolic quadrics. Section 3 deals with the two types of hexagon’s embeddings. We first discuss the principal difference between them making use of spreads of planes. Then we show how the intersections of a skew-embedded hexagon with various doilies help us to better understand the fact that the complement of this hexagon is a contextual configuration. Section 4, the core section of the paper, first highlights the facts that the degree of contextuality of the configuration comprising all 315 line-contexts of $\mathcal{W}(5,2)$ is equal to $63$ and that the corresponding 63 unsatisfiable contexts of this configuration are the 63 lines of a copy of the split Cayley hexagon of order two that is embedded classically into $\mathcal{W}(5,2)$. The first property is then substantiated by a chain of group-theoretic and algebro-geometric arguments. After introducing a line-layered decomposition of the hexagon and describing how this decomposition helps us to quickly alternate between the two embeddings, we illustrate on several examples another crucial finding, namely that contextuality properties of most prominent families of contextual three-qubit configurations having three-element contexts, entailing doilies, elliptic quadrics and hyperbolic quadrics, are fully described in terms of their intersection with properly-selected classically-embedded copies of the hexagon. As a substantiation of our findings, Section 5 outlines the results of testing a specific contextual inequality introduced by Cabello on the IBM Quantum Experience by employing an elliptic quadric and the complement of a particular copy of skew-embedded hexagon of $\mathcal{W}(5,2)$. Our procedure follows and improves that of [holweck2021testing], where the contextuality of $\mathcal{W}(5,2)$ as a whole was already successfully tested. Finally, Section 6 summarises the main achievements and makes some proposals of how to tackle the next case, that of four qubits, in a similar unifying way.
@@ 2 Contextual configurations in symplectic polar spaces, degree of contextuality
Over the past $15$ years symplectic geometry over the two-elements field has been investigated in the context of quantum information as it encodes the commutation relations between the elements of the $n$-qubit Pauli group [saniga2007multiple, havlicek2009factor]. Recall that the Pauli matrices $X,Y,Z$ and the identity operator $I$ can be given in terms of $X$ and $Z$ as follows (phase omitted): $X=Z^{0}X,Y=Z^{1}X^{1},Z=Z^{1}X^{0}$ and $I=Z^{0}X^{0}$. Consider the map:
$\begin{array}[]{ccc}\mathcal{P}_{n}/\{\pm I^{\otimes n},\pm iI^{\otimes n}\}&\to&\mathbb{F}_{2}^{2n}\\ (Z^{a_{1}}X^{b_{1}})\otimes(Z^{a_{2}}X^{b_{2}})\otimes\dots\otimes(Z^{a_{n}}X^{b_{n}})&\mapsto&(a_{1},a_{2},\dots,a_{n},b_{1},b_{2},\dots,b_{n}).\end{array}$
This map associates bijectively to any $n$-qubit Pauli matrix (up to a phase) a unique vector of $\mathbb{F}_{2}^{2n}$. Thinking projectively, one can associate to any (phase disregarded) non-trivial operator of $\mathcal{P}_{n}/\{\pm I^{\otimes n},\pm iI^{\otimes n}\}$ a unique point of PG$(2n-1,2)$, the projective space over the two-elements field.
Let us equip the projective space PG$(2n-1,2)$ with the symplectic form defined as
$\sigma(x,y)=x_{1}y_{n+1}+x_{2}y_{n+2}+\dots x_{n}y_{2n}+x_{n+1}y_{1}+x_{n+2}y_{2}+\dots+x_{2n}y_{n}.$
Then the space $\mathcal{W}(2n-1,2)$ of all totally isotropic subspaces of PG$(2n-1,2)$ with respect to $\sigma$ is called the symplectic polar space of rank $n$ and order $2$. This space encodes the commutation relations of $\mathcal{P}_{n}$ in the sense that one can check by a straightforward calculation that any pair of isotropic elements $x,y\in\mathcal{W}(2n-1,2)$ corresponds to two commuting observables $O_{x}$ and $O_{y}$ in $\mathcal{P}_{n}$, see e. g. [havlicek2009factor].
A large number of contextual configurations can thus be advantageously identified with subgeometries of $\mathcal{W}(2n-1,2)$. Indeed, as contexts are sets of mutually commuting observables whose products are, up to a sign, equal to the identity operator, then all lines, planes, and more generally linear subspaces of $\mathcal{W}(2n-1,2)$ can be chosen as contexts. For example, the Fano plane depicted in Figure 2 is a linear subspace of the largest dimension in $\mathcal{W}(5,2)$. The most prominent subgeometries of $\mathcal{W}(2n-1,2)$ are quadrics. A hyperbolic quadric $\mathcal{Q}^{+}(2n-1,2)$, $n\geq 1$, is a subgeometry whose equation can be brought to the following standard form
@@
points and there are
@@
points and $\mathcal{W}(2n-1,2)$ features
@@
copies of them.
When it comes to $\mathcal{W}(5,2)$, here we find a particularly remarkable kind of subgeometry, namely the split Cayley hexagon of order two, $\mathcal{H}$, which will be the central object of our paper. To introduce its definition, let us consider the parabolic quadric $\mathcal{Q}$ of PG$(6,2)$ defined by the following quadratic form
@@
Then (see, e. g., [mal]) the 63 points of $\mathcal{Q}$ and those 63 lines of $\mathcal{Q}$ whose Grassmannian coordinates $p_{ij}\equiv x_{i}y_{j}-x_{j}y_{i}$ satisfy the following equations
@@
define a point-line incidence structure isomorphic to $\mathcal{H}$. In fact, this description of $\mathcal{H}$ is a particular type of the embedding of $\mathcal{H}$ into $\mathcal{Q}$, called classical. There exists, however, another type of embedding of $\mathcal{H}$ into $\mathcal{Q}$, discovered by Coolsaet [cool] and referred by him to as skew, which is furnished by the coordinate map
@@
All the points of a classically embedded $\mathcal{H}$ are on the same footing. This is, however, not the case with a skew embedded $\mathcal{H}$, where three points that behave ‘classically’ have a special footing; these points lie on a line, called the axis by Coolsaet. As we work over $\mathbb{F}_{2}$, we can project $\mathcal{Q}$ into $\mathcal{W}(5,2)$ using the inverse operation to
@@
to obtain the two (types of) embeddings of $\mathcal{H}$ in $\mathcal{W}(5,2)$.
The notion of degree of contextuality was introduced in [DHGMS22] in order to better understand contextual configurations. The degree of contextuality of a configuration of observables is the minimal number of contexts that cannot be satisfied by any NCHV model. This degree is the Hamming distance between the image of the incidence matrix of the point-line incidence structure defining the configuration and the vector encoding the signs of the contexts [DHGMS22]. Let $(O,C)$ be a quantum contextual configuration with $p=|O|$ observables $O=\{M_{1},\ldots,M_{p}\}$ and $l=|C|$ contexts $C=\{c_{1},\ldots,c_{l}\}$. Its incidence matrix $A\in\mathbb{F}_{2}^{l\times p}$ is defined by $A_{i,j}=1$ if the $i$-th context $c_{i}$ contains the $j$-th observable $M_{j}$. Otherwise, $A_{i,j}=0$. Its valuation vector $E\in\mathbb{F}_{2}^{l}$ is defined by $E_{i}=0$ if $e(c_{i})=1$ and $E_{i}=1$ if $e(c_{i})=-1$, where $e$ is the context valuation of $(O,C)$ i. e. $e(c)=1$ if the context $c$ is positive and $e(c)=-1$ if it is negative. Then the degree $d$ of contextuality of $(O,C)$ is defined as follows
@@
where $d_{H}$ is the Hamming distance on the vector space $\mathbb{F}_{2}^{l}$. For example, if one considers the Magic Mermin pentagram in Figure 3, its degree of contextuality is $1$.
In the two-qubit case, the Mermin-Peres magic square and the configuration comprising all 15 three-element contexts of two-qubit Pauli matrices, aka $\mathcal{W}(3,2)$, are both contextual with their degree of contextuality being, respectively, $1$ and $3$. Figure 4 illustrates how these configurations can be parametrized by two-qubit observables. The lines/contexts with a negative sign are doubled. It is clear that assigning $+1$ to each node provides a classical model that satisfies all conditions except those imposed by the negative lines. The fact that the degree of contextuality is $1$ and $3$ indicates that one cannot do better with any NCHV model. Note that the grid is a subgeometry of $\mathcal{W}(3,2)$, being isomorphic to $\mathcal{Q}^{+}(3,2)$; it can be shown that there are in fact ten copies of the grid, i. e. the triangle-free configuration with nine points and six lines, with three points on a line and two lines through a point, lying in $\mathcal{W}(3,2)$. These two types of contextual two-qubit configurations have two remarkable properties that, as we will see, are absent in the three-qubit (and very likely in any higher rank) case. The first one is the fact that the unsatisfiable contexts do not cover all the observables of the configuration. In both cases we have six non-covered observables. In a grid they lie on two disjoint lines, each being also disjoint from the line representing the single unsatisfiable context. In the doily they form a pair of tricentric triads, namely that pair that is the complement to the grid accommodating the three unsatisfiable constraints. The other notable fact is that, save for a single copy of the grid that features three negative contexts, the corresponding unsatisfiable contexts can be identified in each configuration with its negative contexts.
The above-discussed two two-qubit observable-based proofs of the KS Theorem can be considered, together with the Mermin pentagram, as geometrical building blocks of contextuality. Indeed, as recently proved by two of us [mg23], once a configuration has been found to be contextual for a particular labeling by Pauli observables, then one can deduce that the same geometric configuration will be contextual and will have the same degree of contextuality no matter what admissible multi-qubit Pauli parametrization is employed. In particular, the fact that the grids and $\mathcal{W}(3,2)$ are contextual implies that $\mathcal{W}(2n-1,2)$ is contextual for all $n\geq 2$ because the symplectic polar space of a given dimension always contains copies of symplectic polar spaces of a smaller dimension.
$IZ$$ZI$$ZZ$$XI$$IX$$XX$$XZ$$ZX$$YY$$IY$$ZI$$ZY$$XY$$YX$$XI$$ZZ$$IZ$$XZ$$YY$$YI$$XX$$ZX$$IX$$YZ$

### v3

# Hexagons govern three-qubit contextuality

@@ Abstract
Split Cayley hexagons of order two are distinguished finite geometries living in the three-qubit symplectic polar space in two different forms, called classical and skew. Although neither of the two yields observable-based contextual configurations of their own, classically-embedded copies are found to fully encode contextuality properties of the most prominent three-qubit contextual configurations in the following sense: for each set of unsatisfiable contexts of such a contextual configuration there exists some classically-embedded hexagon sharing with the configuration exactly this set of contexts and nothing else. We demonstrate this fascinating property first on the configuration comprising all 315 contexts of the space and then on doilies, both types of quadrics as well as on complements of skew-embedded hexagons. In connection with the last-mentioned case and elliptic quadrics we also conducted some experimental tests on a Noisy Intermediate Scale Quantum (NISQ) computer to substantiate our theoretical findings.
@@ 1 Introduction
Finite geometry is a branch of mathematics that deals with geometries made of a finite number of points, lines and/or linear spaces of higher dimensions. This perspective of working with spaces that contain only a finite number of geometric elements is rather counter-intuitive and far from the intuition provided, for example, by Euclidean geometry as it lacks concepts like smoothness, differentiability, distance etc. Gino Fano [12] was one of the first geometers to formalize this idea and his name is now associated with the smallest finite projective plane, i. e. the Fano plane (Figure 1).
Figure 1: The Fano plane depicted in its standard rendering. This self-dual geometry comprises seven points (bullets) and seven lines (six straight segments and a circle), with three points per line and, dually, three lines through a point, being equivalent to the projective plane over the two-element field $\mathbb{F}_{2}=\{0,1\}$.
Over the past 20 years (see, for example, [36, 15, 35, 24, 26, 41] and/or [46, 47] for a slightly different, more heuristic approach), finite geometry has been introduced into the field of quantum information and mathematical physics to mainly model and analyse the commutation relations within the $n$-qubit Pauli group, $\mathcal{P}_{n}$, defined by
@@
where $X$, $Y$ and $Z$ are the famous Pauli matrices, $I$ is the associated $2\times 2$ identity matrix and
$A_{1}A_{2}\dots A_{n}\equiv A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}.$ (2)
For example, for $n=3$, if one ignores the global phase of each operator, a maximum set of mutually commuting three-qubit Pauli operators (disregarding the identity) forms a Fano plane, as shown with the example provided by Figure 2. Two points are collinear if and only if the matrix product of the observables labelling them commutes.
@@
In what follows, a positive or negative (quantum) context will be a set of mutually commuting observables such that their product is $+I^{\otimes n}$ or $-I^{\otimes n}$, respectively, where $I^{\otimes n}\equiv I_{(1)}\otimes I_{(2)}\otimes\dots\otimes I_{(n)}$ is the $n$-fold tensor product of $2\times 2$ identity matrices. For instance, the seven observables of Figure 2 form a (negative) three-qubit context. A contextual configuration will be an arrangement of observables made of contexts such that there is no Non-Contextual Hidden Variable (NCHV) model that can reproduce the outcomes predicted by the rules of Quantum Mechanics (QM).
@@
Let us consider, for example, the configuration portrayed in Figure 3, known as the Mermin pentagram. Each node of the configuration is a three-qubit Pauli operator whose eigenvalues are $+1$ and $-1$. Each line of the configuration represents a context. The doubled line is the unique context with a negative sign, i. e. the product of the observables on this context is $-I^{\otimes 3}$. David Mermin introduced this configuration in [27] as an alternative proof of the famous Kochen-Specker (KS) Theorem  (see, for example, [21, 3, 40, 4, 6]). Recall that the Kochen-Specker Theorem is a no-go result that proves the non-existence of an NCHV model, i. e. it proves that a Hidden Variable (HV) model that would reproduce the outcomes predicted by QM has to be context-dependent. Consider our Mermin pentagram. An HV model that reproduces predictions of QM for this set of operators should satisfy the following constraints:
@@
Observable-based proof of the KS Theorem can already be built from two-qubit Pauli observables, like in the Mermin-Peres magic square, see Section 2 and [31, 27]. For more information on the two-qubit case, the reader is referred to consult [10, 19], as well as [6] for a recent survey on contextuality.
@@
In this paper we will focus solely on the three-qubit case ($n=3$) to demonstrate in detail how a specific arrangement of three-qubit contexts isomorphic to the smallest split Cayley hexagon underpins contextuality properties of a whole class of distinguished aggregates of three-qubit observables. This hexagon made its debut in physics as early as 2008 in [25], where the authors employed its particular three-qubit subgeometry to model the $E_{7}$-symmetric black-hole entropy formula in string theory. Later [32], it was pointed out that the order of the automorphism group of the hexagon, 12 096, coincides with the total number of three-qubit Mermin pentagrams. At about the same time [38], it was already noticed that some contextual three-qubit configurations can be uniquely extended to geometric hyperplanes of the hexagon.
@@
It was however only very recently [18] that the fact that the hexagon embeds in the three-qubit space in two different ways has been properly taken into account and recognized as having deep physical meaning, leading to our present study.
@@
The paper is organized as follows. In Section 2 we introduce the notion of the symplectic polar space of rank $n$ and order two, $\mathcal{W}(2n-1,2)$, which is the geometric framework for the commutation relations within the $n$-qubit generalized Pauli group. Next, we define the most prominent subgeometries of this space, quadrics, and show that the particular space $\mathcal{W}(5,2)$ for three-qubits is endowed with another remarkable kind of subgeometry, namely the split Cayley hexagon of order two, which occurs in this space in two non-isomorphic embeddings. Then we introduce the notion of the degree of contextuality of a quantum configuration, that measures how far a contextual configuration of observables is from supporting an NCHV model, and explicitly illustrate this notion on the example of the smallest symplectic space $\mathcal{W}(3,2)$, the doily, and one of its hyperbolic quadrics. Section 3 deals with the two types of hexagon’s embeddings. We first discuss the principal difference between them making use of sets of nine mutually disjoint planes of $\mathcal{W}(5,2)$. Then we show how the intersections of a skew-embedded hexagon with various doilies (i. e. $\mathcal{W}(3,2)$’s) help us to better understand the fact that the complement of this hexagon is a contextual configuration. Section 4, the core section of the paper, first highlights the facts that the degree of contextuality of the configuration comprising all 315 line-contexts of $\mathcal{W}(5,2)$ is equal to $63$ and that the corresponding 63 unsatisfiable contexts of this configuration are the 63 lines of a copy of the split Cayley hexagon of order two that is embedded classically into $\mathcal{W}(5,2)$. The first property is then substantiated by a chain of group-theoretic and algebro-geometric arguments. After introducing a line-layered decomposition of the hexagon and describing how this decomposition helps us to quickly alternate between the two embeddings, we illustrate on several examples our most crucial finding, namely that contextuality properties of most prominent families of contextual three-qubit configurations having three-element contexts, entailing doilies, elliptic quadrics and hyperbolic quadrics, are fully described in terms of their intersection with properly-selected classically-embedded copies of the hexagon. As a substantiation of our findings, Section 5 outlines the results of testing a specific contextual inequality introduced by Cabello on the IBM Quantum Experience by employing an elliptic quadric and the complement of a particular copy of skew-embedded hexagon of $\mathcal{W}(5,2)$. Our procedure follows and improves that of [17], where the contextuality of $\mathcal{W}(5,2)$ as a whole was already successfully tested. Finally, Section 6 summarizes the main achievements, makes some proposals of how to tackle the next case, that of four qubits, in a similar unifying way and briefly addresses an intriguing formal analogy between the two inequivalent embeddings of the smallest split Cayley hexagon into the three-qubit $\mathcal{W}(5,2)$ and two inequivalent kinds of genuine tripartite entanglement.
@@ 2 Contextual configurations in symplectic polar spaces, degree of contextuality
Over the past $15$ years symplectic geometry over the two-element field has been investigated in the context of quantum information as it encodes the commutation relations between the elements of the $n$-qubit Pauli group [37, 14]. Recall that the Pauli matrices $X,Y,Z$ and the identity operator $I$ can be given in terms of $X$ and $Z$ as follows (phase omitted): $X=Z^{0}.X,Y=Z^{1}.X^{1},Z=Z^{1}.X^{0}$ and $I=Z^{0}.X^{0}$. Consider the map:
$\begin{array}[]{ccc}\mathcal{P}_{n}/\{\pm I^{\otimes n},\pm iI^{\otimes n}\}&\to&\mathbb{F}_{2}^{2n}\\ (Z^{a_{1}}.X^{b_{1}})\otimes(Z^{a_{2}}.X^{b_{2}})\otimes\dots\otimes(Z^{a_{n}}.X^{b_{n}})&\mapsto&(a_{1},a_{2},\dots,a_{n},b_{1},b_{2},\dots,b_{n}).\end{array}$ (3)
This map associates bijectively to any $n$-qubit Pauli matrix (up to a phase) a unique vector of $\mathbb{F}_{2}^{2n}$. Thinking projectively, one can associate to any (phase disregarded) non-trivial operator of $\mathcal{P}_{n}/\{\pm I^{\otimes n},\pm iI^{\otimes n}\}$ a unique point of PG$(2n-1,2)$, the $(2n-1)$-dimensional projective space over the two-element field [^1] Let $V$ be a ($d+1$)-dimensional vector space over $\mathbb{F}_{2}$. The projective space PG$(d,2)$ is the geometry whose points, lines, planes, etc. are the vector subspaces of $V$ of respective dimensions 1, 2, 3, etc.. Its points can be represented by ($d+1$)-tuples of the form $(x_{1},x_{2},x_{3},\ldots,x_{d+1})$ where $x_{i}\in\mathbb{F}_{2}$, disregarding the trivial $(0,0,0,\ldots,0)$-tuple..
Let us equip the projective space PG$(2n-1,2)$ with the symplectic form defined as (note that $+1=-1$ over $\mathbb{F}_{2}$)
$\sigma(x,y)=x_{1}y_{n+1}+x_{2}y_{n+2}+\dots+x_{n}y_{2n}+x_{n+1}y_{1}+x_{n+2}y_{2}+\dots+x_{2n}y_{n}$ (4)
and call a subspace of PG$(2n-1,2)$ totally isotropic if this form vanishes identically on it, i. e. $\sigma(x,y)=0$ for any two distinct points $x$ and $y$ of the subspace. Then the space $\mathcal{W}(2n-1,2)$ of all totally isotropic subspaces of PG$(2n-1,2)$ with respect to $\sigma$ is called the symplectic polar space of rank $n$ and order $2$. This space encodes the commutation relations of $\mathcal{P}_{n}$ in the sense that one can check by a straightforward calculation that any pair of collinear points $x,y\in\mathcal{W}(2n-1,2)$ corresponds to two commuting observables $O_{x}$ and $O_{y}$ in $\mathcal{P}_{n}$, see e. g. [14]. In what follows, it will be implicitly assumed that all the points of such multi-qubit $\mathcal{W}(2n-1,2)$ are labelled solely by canonical observables of $\mathcal{P}_{n}$, i. e. by those observables whose phase $s$ equals $1$ (see eq. (1)). Moreover, when referring to a subspace of $\mathcal{W}(2n-1,2)$ we will always have in mind also the associated set of pairwise commuting observables, and vice versa.
A large number of contextual configurations can thus be advantageously identified with subgeometries of $\mathcal{W}(2n-1,2)$. Indeed, as contexts are sets of mutually commuting observables whose products are, up to a sign, equal to the identity operator, then all lines, planes, and more generally linear subspaces of $\mathcal{W}(2n-1,2)$ can be chosen as contexts. For example, the Fano plane depicted in Figure 2 is a linear subspace of the largest dimension in $\mathcal{W}(5,2)$. The most prominent subgeometries of $\mathcal{W}(2n-1,2)$ are quadrics, which occur in two different forms. A hyperbolic quadric $\mathcal{Q}^{+}(2n-1,2)$, $n\geq 1$, is a subgeometry whose equation can be brought to the following standard form
@@
points,
$|\mathcal{Q}^{+}|_{l}=\frac{1}{3}(2^{n}-1)(2^{n-2}+1)(2^{2(n-1)}-1)$ (7)
lines and there are
@@
points,
$|\mathcal{Q}^{-}|_{l}=\frac{1}{3}(2^{n}+1)(2^{n-2}-1)(2^{2(n-1)}-1)$ (11)
lines and $\mathcal{W}(2n-1,2)$ features
@@
[…]
When it comes to $\mathcal{W}(5,2)$, here we find a particularly remarkable kind of subgeometry, namely the split Cayley hexagon of order two, $\mathcal{H}$, which will be the central object of our paper. It is a finite geometry having 63 points and 63 lines, with three points on a line and three lines through a point, such that its smallest polygons are hexagons (hence its name). To introduce its abstract definition, let us consider the parabolic quadric $\mathcal{Q}$ in the six-dimensional projective space PG$(6,2)$ defined by the following quadratic form
@@
Then (see, e. g., [43]) the 63 points of $\mathcal{Q}$ and those 63 lines of $\mathcal{Q}$ whose Grassmannian coordinates $p_{ij}\equiv x_{i}y_{j}-x_{j}y_{i}$ satisfy the following equations
@@
define a point-line incidence structure isomorphic to $\mathcal{H}$. In fact, this description of $\mathcal{H}$ is a particular type of the embedding of $\mathcal{H}$ into $\mathcal{Q}$, called classical. There exists, however, another type of embedding of $\mathcal{H}$ into $\mathcal{Q}$, discovered by Coolsaet [9] and referred by him to as skew, which is furnished by the coordinate map
@@
All the points of a classically-embedded $\mathcal{H}$ are on the same footing. This is, however, not the case with a skew-embedded $\mathcal{H}$, where three points that behave ‘classically’ have a special footing; these points lie on a line, called the axis by Coolsaet. As we work over $\mathbb{F}_{2}$, we can project $\mathcal{Q}$ into $\mathcal{W}(5,2)$ using the inverse operation to
@@
to obtain the two (types of) embeddings of $\mathcal{H}$ in $\mathcal{W}(5,2)$. Similarly to quadrics, also in this case we can completely avoid working with the above-given abstract definitions. There already exists [39, 33] very appealing, and also highly symmetric, graphical visualisations/drawings of the smallest split Cayley hexagon that illustrate all its essential geometric properties and exhibit all its 63 points and 63 lines in a particularly handy way to work with. So, the exposition of our ideas in Sections 3 and 4 will almost exclusively rest on such drawings, where the points of the hexagon are labeled by canonical three-qubit observables in conformity with the mapping defined by (3). As each hexagon contains all 63 points of $\mathcal{W}(5,2)$, the embedding of a given copy of the hexagon into this space is encoded solely in the way how its points are ‘arranged’ into lines of $\mathcal{W}(5,2)$, which in a diagrammatic language translates to a specific way of how the points of the underlying figure are labelled by the 63 three-qubit observables. This also means that one can pass from one copy of the hexagon to the other by a simple permutation of these three-qubit labels on the figure in question, the only constraint to be secured during such a procedure is that the three observables associated with any line mutually commute and their (ordinary) product is proportional to $I^{\otimes 3}$. We will further see that a special kind of permutation that transforms a classically-embedded hexagon to a skew-embedded one, and vice versa, is intricately related to a line-related layering of the hexagon, this layering being also essential in understanding the (pivotal) role played by the axis of a skew-embedded copy under such a transformation. Also, we shall explicitly illustrate the intersection of selected representatives of the two kinds of quadrics with a classically-embedded hexagon to demonstrate its most baffling contextuality-related role. And all that in a way to be also accessible to the reader having only a very limited background in finite geometry.
The notion of degree of contextuality was introduced in [10] in order to better understand contextual configurations. The degree of contextuality of a configuration of observables is the minimal number of contexts that cannot be satisfied by any NCHV model. This degree is the Hamming distance between the image of the incidence matrix of the point-line incidence structure defining the configuration and the vector encoding the signs of the contexts [10]. Let $(O,C)$ be a quantum contextual configuration with $p=|O|$ observables $O=\{M_{1},\ldots,M_{p}\}$ and $l=|C|$ contexts $C=\{c_{1},\ldots,c_{l}\}$. Its incidence matrix $A\in\mathbb{F}_{2}^{l\times p}$ is defined by $A_{i,j}=1$ if the $i$-th context $c_{i}$ contains the $j$-th observable $M_{j}$. Otherwise, $A_{i,j}=0$. Its valuation vector $E\in\mathbb{F}_{2}^{l}$ is defined by $E_{i}=0$ if $e(c_{i})=1$ and $E_{i}=1$ if $e(c_{i})=-1$, where $e$ is the context valuation of $(O,C)$ i. e. $e(c)=1$ if the context $c$ is positive and $e(c)=-1$ if it is negative. Then the degree $d$ of contextuality of $(O,C)$ is defined as follows
@@
where $d_{H}$ is the Hamming distance on the vector space $\mathbb{F}_{2}^{l}$. In order to find this degree we proceed as follows. Given a quantum configuration ${K}=(O,C)$, one associates with it a configuration $\widetilde{{K}}$ that is geometrically identical with ${K}$, but has all its observables replaced by $+1$’s and $-1$’s and the sign of each context $c\in C$ replaced by the product of these $+1$’s and $-1$’s over its members. Then we perform an exhaustive, computer-aided search – which basically consists of reshuffling and/or swapping these numbers following specific algorithms described in detail in [29, 30] – to find such a $\widetilde{{K}}$ that has the maximum possible numbers of lines having the same sign as the corresponding contexts in ${K}$. The degree of contextuality of ${K}$ is then equal to the number of remaining lines of $\widetilde{{K}}$, i. e. those lines each of which has different parity than its corresponding context in ${K}$; these contexts of ${K}$ that have different signs than the corresponding lines of $\widetilde{{K}}$ are called unsatisfiable.
$IZ$$ZI$$ZZ$$XI$$IX$$XX$$XZ$$ZX$$YY$
$IY$$ZI$$ZY$$XY$$YX$$XI$$ZZ$$IZ$$XZ$$YY$$YI$$XX$$ZX$$IX$$YZ$